### v2

# Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions

Abstract.
In recent years, simultaneous learning of multiple dense prediction tasks with partially annotated label data has emerged as an important research area. Previous works primarily focus on leveraging cross-task relations or conducting adversarial training for extra regularization, which achieve promising performance improvements, while still suffering from the lack of direct pixel-wise supervision and extra training of heavy mapping networks. To effectively tackle this challenge, we propose a novel approach to optimize a set of compact learnable hierarchical task tokens, including global and fine-grained ones, to discover consistent pixel-wise supervision signals in both feature and prediction levels. Specifically, the global task tokens are designed for effective cross-task feature interactions in a global context. Then, a group of fine-grained task-specific spatial tokens for each task is learned from the corresponding global task tokens. It is embedded to have dense interactions with each task-specific feature map. The learned global and local fine-grained task tokens are further used to discover pseudo task-specific dense labels at different levels of granularity, and they can be utilized to directly supervise the learning of the multi-task dense prediction framework. Extensive experimental results on challenging NYUD-v2, Cityscapes, and PASCAL Context datasets demonstrate significant improvements over existing state-of-the-art methods for partially annotated multi-task dense prediction.
Multi-task Partially Annotated Dense Prediction, Multi-task Learning, Label Discovery
submissionid: 5591copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746027.3755727conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision tasks
1. Introduction
With the rapid development of supervised learning with deep neural networks, various pixel-wise dense prediction tasks with highly complementary properties such as semantic segmentation and depth estimation have achieved great success in multi-task learning (MTL) in recent years (Misra et al., 2016; Vandenhende et al., 2020; Xu et al., 2018; Ye and Xu, 2022; Zhang et al., 2018). Researchers pursue the learning of them simultaneously in a unified framework, which can effectively model cross-task correlations and achieve superior results in terms of model training costs and performances.
However, in real-world scenarios, obtaining pixel-level annotations is prohibitively expensive, especially when dealing with a set of distinct dense prediction tasks. Each image has to be annotated with pixel labels for all the tasks. Thus, existing works have delved into the problem of multi-task learning with only partially annotated dense labels (Li et al., 2022; Luo et al., 2021; Wang et al., 2022c; Zamir et al., 2020; Zeng et al., 2019; Nishi et al., 2024; Ye and Xu, 2024). Specifically, as illustrated in Fig. 1 (a), given an input image, for $T$ dense prediction tasks, the task labels are provided partially, i.e. for at least one task and at most $T-1$ tasks. Learning a multi-task model under this setting is particularly challenging since every input image lacks some of the task supervision signals, and the performance typically drops significantly if compared to the same model trained with full task label supervisions (Li et al., 2022).
Figure 1. (a) Illustration of partially annotated multi-task dense prediction setting. Each input image only has partial task labels from all the tasks. (b) Illustration of the learning of Hierarchical Task Tokens, including global task tokens and fine-grained task tokens, by conducting feature-token interactions in feature and prediction spaces separately. The well-learned hierarchical task tokens can achieve both feature supervision discovery and task label discovery.
Previous works have been focusing on excavating cross-task relations by training heavy extra mapping networks (Li et al., 2022; Nishi et al., 2024), however, simply applying regularization in compact latent spaces fails to address the lack of dense pixel supervision and limits the performance. On the contrary, directly discovering pseudo task labels in prediction spaces can alleviate this problem to a certain extent, while still suffering from the following two severe limitations: (i) Simply discovering task labels in the prediction spaces separately ignores the highly relevant task relations, thus leading to trivial performance under multitask scenarios. (ii) Solely discovering labels in prediction spaces cannot take advantage of the abundant task representations in feature space. For each specific task, the distributions of the task features and predictions should be consistent. Since the encoder parameters are shared among tasks, thus the produced rich task-generic features are beneficial for building cross-task relations and discovering supervision signals effectively. Therefore, it is critically important to involve hierarchical cross-task representations from feature space to prediction space to consistently boost the task label discovery process.
To effectively tackle the aforementioned challenges, we propose a novel approach that performs task label discovery from both the feature and prediction spaces via an effective design of learnable Hierarchical Task Tokens (HiTTs). HiTTs are sets of compact parameters that are learned in a hierarchical manner to model global inter-task relationships and local fine-grained intra-task relationships, which allows for discovering pixel-wise task pseudo labels straightforwardly in both feature and prediction spaces consistently. More specifically, as depicted in Fig. 1 (b), we apply HiTTs during the multi-task decoding stage and jointly optimize them with the multi-task learning network. The HiTTs consist of two hierarchies. The first hierarchy is a set of global task tokens. The global task tokens are randomly initialized and can perform cross-task feature-token interactions with different task feature maps based on self-attention. These learned task tokens can be used to discover feature-level pseudo supervision by selecting highly activated pixel features correlated to each task. The second hierarchy is the fine-grained task tokens. These tokens are directly derived from the global task tokens with learnable projection layers to inherent beneficial task representations. They are subsequently utilized to perform interactions within each task-specific feature map at a finer granularity. As the fine-grained task tokens can learn pixel-to-pixel correlation with each task feature map, they thus benefit the discovery of dense spatial labels for each task. Compared with naive pseudo labeling process (Lee et al., 2013), the HiTTs can bridge effective information from supervised tasks so as to encourage highly confident predictions on unsupervised tasks. We learn both hierarchies of tokens simultaneously in an end-to-end manner incorporating the multi-task baseline network, and exploit both levels of supervision signals discovered from the two hierarchies, for optimizing multi-task dense predictions on partially annotated datasets.
In summary, the contribution of this work is three-fold:
  • Instead of discovering cross-task regularization by extra heavy mapping networks, we propose to utilize cross-task relations for high-quality task label discovery, which serves as pixel-level dense pseudo supervision under the multi-task partially supervised setting.
  • We propose a novel design of Hierarchical Task Tokens (HiTTs), which can learn hierarchical multi-task representations for high-quality pseudo label discovery consistently in both the feature and the prediction levels.
  • Our proposed method significantly outperforms existing state-of-the-art competitors on multi-task partially annotated benchmarks, including PASCAL-Context, NYUD-v2 and Cityscapes, and demonstrates clear effectiveness on challenging dense prediction tasks with limited annotations, including segmentation, depth estimation, normal estimation and edge detection, etc. Code is released at https://github.com/Evergreen0929/EEMTL.
Figure 2. Illustration of our method. HiTTs consist of both global and fine-grained task tokens which learn discriminative task representations by conducting feature-token interactions with attentions in corresponding multi-task decoding stages. The global task tokens $\boldsymbol{\theta_{i}}$ discover feature-level pseudo supervision $\mathcal{L}_{f}$, while the fine-grained task tokens $\boldsymbol{\varphi_{i}}$ inherit the knowledge from global task tokens and directly discover pixel labels for supervision $\mathcal{L}_{p}$. The supervision from ground-truth label is denoted as $\mathcal{L}_{s}$.
2. Related Work
Multi-task Dense Prediction. Dense prediction tasks aim to produce pixel-wise predictions for each image. Common tasks including semantic segmentation, depth estimation, and surface normal estimation exhibit high cross-task correlations. For instance, depth discontinuity is usually aligned with semantic boundaries (Vandenhende et al., 2020), and surface normal distributions are aligned with spatial derivatives of the depth maps (Lu et al., 2021). Thus, a number of works have been focusing on multi-task dense predictions (Gao et al., 2019; Liu et al., 2019; Misra et al., 2016; Vandenhende et al., 2020; Xu et al., 2018; Yang et al., 2023b; Ye and Xu, 2022, 2023a, 2023b; Zhang et al., 2023, 2018, 2019, 2025; Tang et al., 2025; Tian et al., 2024; Lu et al., 2024; Cao et al., [n. d.]; Yang et al., [n. d.]; Chavhan et al., 2025). They leverage parameters sharing to conduct cross-task interactions by effective attention mechanisms for task feature selection (Liu et al., 2019), and multi-modal distillation (Xu et al., 2018), multi-scale cross-task interactions (Vandenhende et al., 2020), global pixel and task interactions (Ye and Xu, 2022) and multi-task bridge features (Zhang et al., 2025). However, these works focus on fully-supervised settings. In contrast, our work addresses the challenge of insufficient supervision signals in each task.
Semi-supervised learning. Obtaining pseudo labels for semi-supervised learning is a popular research direction, with several deep learning works published on the topic (Iscen et al., 2019; Lee et al., 2013; Li et al., 2019; Shi et al., 2018; Sohn et al., 2020; Tarvainen and Valpola, 2017; Xie et al., 2020; Zou et al., 2018, 2019). Among them, (Lee et al., 2013) aims at picking up the class which has the maximum predicted confidence. The graph-based label propagation method (Iscen et al., 2019) is also used to infer pseudo-labels for unlabeled data. (Shi et al., 2018) provides a confidence level for each unlabeled sample to reduce influences from outliers and uncertain samples, and uses MMF regularization at feature levels to make images with the same label close to each other in the feature space. (Xie et al., 2020) uses accurate pseudo labels produced by the teacher model on clean unlabeled data to train the student model with noise injected. For semi-supervised dense prediction tasks (Pieropan et al., 2022; Yang et al., 2023a; Hoyer et al., 2024; Mai et al., 2024; Ran et al., 2024; Yang et al., 2025), such as semantic segmentation, several works focus on assigning pixel-wise pseudo annotations from high-confidence predictions (Li et al., 2019; Zou et al., 2018, 2019). However, these works target single-task learning setups. Despite pseudo labeling, discovering consistency for regularization is also a popular direction for unlabeled data (Li et al., 2022; Luo et al., 2021; Tang and Jia, 2022; Zeng et al., 2019). (Li et al., 2022; Luo et al., 2021; Zeng et al., 2019) focus on building cross-task consistency, while  (Tang and Jia, 2022) uses image-level feature similarities to find important samples for semi-supervised learning. Differently, our work targets pixel-level task supervision discovery by hierarchical task tokens containing multi-level multi-task representations for partially annotated dense predictions.
Multi-task Partially Supervised Learning. As discussed in the introduction, obtaining pixel-level annotations for every task on images is prohibitively expensive. Therefore, some recent works focus on partially annotated settings for multi-task learning (Imran et al., 2020; Li et al., 2022; Liu et al., 2007; Lu et al., 2021; Luo et al., 2021; Wang et al., 2022c; Ye and Xu, 2024; Zamir et al., 2020; Zeng et al., 2019; Nishi et al., 2024). Since directly recovering labels from other tasks is an ill-posed problem (Li et al., 2022), enforcing consistency among tasks is usually adopted. For instance, constructing a common feature space to align predictions and impose regularization (Li et al., 2022), and leveraging intrinsic connections of different task pairs between predictions of different tasks on unlabeled data in a mediator dataset, when jointly learning multiple models (Lu et al., 2021). Adversarial training is also adopted to align the distributions between labeled and unlabeled data by discriminators (Wang et al., 2022c), and multi-task denoising diffusion is adopted (Ye and Xu, 2024) to address the issue of noise in initial prediction maps. To the best of our knowledge, our hierarchical task tokens for both pseudo feature supervision and task label discovery are a novel exploration of the problem, and show a clear difference from existing works.
3. Proposed Method
Our proposed approach for learning Hierarchical Task Tokens (HiTTs) primarily comprises two stages, i.e., the Global Token Learning and the Fine-grained Token Learning. The overall structure of HiTTs is depicted in Fig. 2. Firstly, in the Global Token Learning stage, the global task tokens produce task features and then learn rich task-level representations by conducting inter- and intra-task attention with all task features. The global tokens are utilized to exploit rich representations in feature space and discover feature-level pseudo supervision. Subsequently, in the fine-grained stage, we project each task feature into fine-grained feature space by simple convolution layers, and derive the fine-grained tokens from the global tokens by Multi-layer Perceptrons (MLPs), to inherit well-learned global task representations and therefore achieve consistent pseudo label discovery. To perform a uniform confidence-based pseudo label discovery for different types of dense prediction tasks, we follow (Brüggemann et al., 2021) to conduct discrete quantization of regression task annotations (e.g. depth estimation and normal estimation), and treat all tasks as pixel-wise classification.
Figure 3. (a) Global Task Token Learning. It mainly contains two stages: Inter- and Intra-Task Attention for thoroughly feature-token cross-task interaction to obtain robust and representative global task tokens, and excavating pseudo feature supervision based on the learned task tokens. (b) Fine-grained Task Token Learning. The projected fine-grained tokens and feature maps (projection is shown in Fig. 4) can be used for discovering high-quality pseudo labels in prediction spaces.
3.1. Global Task Token Learning
In the global task token learning stage, we target learning global tokens representing the distributions of each task, which are further used for pseudo feature supervision discovery. The learning process is mainly achieved by inter- and intra-task attention among tokens and features, in order to exploit beneficial multi-task representations for token learning and feature supervision discovery.
Given an RGB input $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a multi-task dense prediction framework firstly produces a task-generic representation $\boldsymbol{F_{s}}\in\mathbb{R}^{C\times h\times w}$ through a shared encoder. Considering we have $T$ tasks, and we target decoding task features $\{\boldsymbol{F_{1}},\boldsymbol{F_{2}},\cdots,\boldsymbol{F_{T}}\}$ from $\boldsymbol{F_{s}}$ as well as learning representative global task tokens $\{\boldsymbol{\theta_{1}},\boldsymbol{\theta_{2}},\cdots,\boldsymbol{\theta_{T}}\},\boldsymbol{\theta_{i}}\in\mathbb{R}^{C}$ for each task. They are randomly initialized learnable vectors serving as additional input tokens for the decoder.
As shown in Fig. 3 (a), the proposed global task token learning process is mainly composed of two stages: i) Inter- and Intra-Task Attention, which aim to thoroughly model feature-token cross-task relations to obtain robust and representative global task tokens. ii) Feature Supervision Discovery, which excavates pseudo feature supervision with confidence maps provided by global task tokens for unsupervised task features.
Firstly, we use each global task token to query the shared feature to obtain each task feature $\boldsymbol{F_{i}}$ accordingly. Then, to conduct inter-task attention, all global task tokens are used to calculate all-task affinities $\boldsymbol{A}\in\mathbb{R}^{T\times T}$ to represent the global task relations. After the $\boldsymbol{A}$ is calculated, it is used to conduct affine combinations of task features and global task tokens respectively. Since for each task feature $\boldsymbol{F_{i}}$, if it is not directly supervised by labels, the feature will be less representative and contain more noise. Thus, conducting affine combinations among all tasks ensures that beneficial discriminative representations from other supervised task features can fertilize the unsupervised ones. Afterward, the updated tokens and features containing cross-task information are fed to the intra-task attention module, where they are rearranged and grouped in each task, and self-attention is applied to each group of task token and feature. The global task tokens will further learn more specific and discriminative task representations during this process, and representative task tokens will in turn enhance the feature quality as well.
Followingly, we discover pseudo feature supervision with the aid of well-learned global task tokens, and this process will be discussed in Sec. 3.3. In addition, for multi-scale backbone features, directly fusing them ignores the various granularity of task representations maintained at different scales. Thus, for multi-scale image backbone, we further propose Multi-scale Global Task Token Learning in order to learn comprehensive multi-scale task relations. The proposed method involves inter-task attention separately at each scale, and then the multi-scale features and tokens are fused before intra-task attention. In this way, the global task tokens gain richer cross-task relations at different scales and are able to maintain stronger representations. The multi-scale global task token learning, inter- and intra-task attention will be illustrated in detail in the supplementary material.
3.2. Fine-grained Task Token Learning
After the global task tokens are learned, we further propose to conduct feature-token interaction at a finer spatial granularity, which takes advantage of various representations in global task tokens and boosts the task label discovery process.
Firstly, as shown in Fig. 4, we jointly project each updated task token $\boldsymbol{\theta_{i}}\in\mathbb{R}^{C}$ and feature $\boldsymbol{F_{i}}\in\mathbb{R}^{C\times h\times w}$ into the prediction space with finer granularity. For features, this can be easily achieved by applying a linear convolution layer, and we denote the fine-grained task features as $\boldsymbol{G_{i}}\in\mathbb{R}^{C_{p}\times H\times W}$, where $C_{p}$ indicates the prediction dimension. For tokens, we denote the projected fine-grained tokens as $\boldsymbol{\varphi_{i}}\in\mathbb{R}^{C_{p}\times C}$, we hope every $1\times C$ vector inside it can represent one category distribution over the spatial dimension. The simplest way is to project each $\boldsymbol{\theta_{i}}$ with a Multi-Layer Perceptron (MLP), which can be described as: $\boldsymbol{\varphi_{i}}=\operatorname{MLP_{i}}(\boldsymbol{\theta_{i}})^{\top},i=1,2,\cdots,T$. However, since there is no direct supervision imposed to distinguish every fine-grain token during this process, the MLPs will tend to degenerate and perform linearly correlated outputs, which prevents the fine-grained tokens from learning discriminative task-specific representations. To alleviate this problem, we propose to use Orthogonal Embeddings (OE) to serve as priors and aid the learning process of fine-grained tokens.
In detail, the vectors in the fine-grained token should be far from each other to represent meaningful and distinguishable task category information, so we use a group of orthogonal basis in $\mathbb{R}^{C_{p}}$ to serve as the embedding for MLP input, denoting as $\boldsymbol{o}\in\mathbb{R}^{C_{p}\times C_{p}}$. These OE are projected into the feature space by linear projections, and then added with the global token before being fed into the MLP. Therefore, with these orthogonal priors, the MLPs can easily keep the distance between the vectors in $\boldsymbol{\varphi_{i}}$ far from each other in the feature space, which makes it able to learn information that distinguishes between task categories, as well as inherit the global representations from $\boldsymbol{\theta_{i}}$.
Subsequently, we exploit the fine-grained task tokens $\boldsymbol{\varphi_{i}}$ for the fine-grained token learning process. Normally, $\boldsymbol{G_{i}}$ is noisy and low-confident on unlabeled data due to the lack of supervision, which leads to inaccurate predictions. We propose to use fine-grained task tokens to encourage $\boldsymbol{G_{i}}$ to produce high-confident logits, and enhance the quality of pseudo labels:
(1) $\boldsymbol{G_{i}^{\prime}}=\operatorname{Conv_{i}^{3\times 3}}\left(\boldsymbol{G_{i}}\odot\operatorname{Softplus}\left(\boldsymbol{\varphi_{i}}\times\boldsymbol{G_{i}}\right)\right),$
where $\operatorname{Softplus}(x)=\log(1+\exp(x))$. Since $\boldsymbol{\varphi_{i}}$ inherits global task representations from $\boldsymbol{\theta_{i}}$, which will perform more robustly on unlabeled data, and aid the production of distinguished logits score in the updated feature $\boldsymbol{G_{i}^{\prime}}$, as well as high-confident final task predictions. Similar to the previous stage, we also conduct pseudo label discovery after the fine-grained tokens are learned, which will be discussed in detail in the next section.
Figure 4. Illustration of Token &\&& Feature Projection and the Fine-grained Task Token Learning. Different tasks have the same structure, and we take one as an example. We project the fine-grained feature $\boldsymbol{G_{i}}$ from the updated task feature $\boldsymbol{F_{i}}$, and derive fine-grained task tokens $\boldsymbol{\varphi_{i}}$ from the updated global task tokens $\boldsymbol{\theta_{i}}$.
3.3. Hierarchical Label Discovery and Multi-task Optimization
For the multi-task partially annotated setting, the training loss on labeled data can be described as:
(2) $\mathcal{L}_{s}=\frac{1}{T}\sum_{i=1}^{T}\left(\alpha_{i}\operatorname{L_{i}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right)\right),$
where where $\operatorname{L_{i}}(\cdot)$ is the loss function for task $i$, and $\alpha_{i}=1$ if task $i$ has ground-truth label, otherwise $\alpha_{i}=0$. $\boldsymbol{\hat{Y_{i}}}$ and ${\boldsymbol{Y_{i}}}$ are task prediction and ground-truth.
We first train the multi-task model with HiTTs jointly with $\operatorname{L_{i}}$ only to achieve convergence on labeled data, then we utilize both hierarchies of tokens to discover feature-level and prediction-level pseudo supervision. As we mentioned in Sec. 3.1, before the updated tokens and features are fed into the next stage, we conduct the feature supervision discovery to excavate feature-level supervision signals for unlabeled tasks. As shown in Fig. 3 (a), we use the updated global task tokens to query each pixel feature on every task and produce confidence mask $\boldsymbol{M_{i}^{f}}=\operatorname{Sigmoid}(\boldsymbol{\theta_{i}}^{\top}\times\boldsymbol{F_{i}})$. Since $\boldsymbol{\theta_{i}}$ is globally learned on all task features, in $\boldsymbol{M_{i}^{f}}$, higher scores indicate that the pixel features have a higher response to task $i$, which should be further used to prove task supervision. Thus we use $\boldsymbol{M_{i}^{f}}$ to serve as a soft confidence mask for pixel-wise feature supervision loss:
(3) $\mathcal{L}_{f}=\frac{1}{T}\sum_{i=1}^{T}\left(\alpha_{i}\operatorname{L_{mse}}\left(\boldsymbol{F_{i}},\boldsymbol{\tilde{F_{i}}}\right)+(1-\alpha_{i})\boldsymbol{M_{i}^{f}}\odot\operatorname{L_{mse}}\left(\boldsymbol{F_{i}},\boldsymbol{\tilde{F_{i}}}\right)\right),$
where $\tilde{\boldsymbol{F_{i}}}$ represents the offline saved features which serve as pseudo supervision signals for unsupervised task features. The $\odot$ represents element-wise multiplication, $\operatorname{L_{mse}}$ is the mean squared error loss for feature distance measurement. For the feature loss on labeled task (first item in Eq 3), we regard all pixel features from $\boldsymbol{\tilde{F_{i}}}$ as valid since they are supervised by ground-truth label, while for unlabeled task (second item in Eq 3), we use $\boldsymbol{M_{i}^{f}}$ encourage high-confidence pixel features and depress low-confidence ones.
Afterward, we also conduct pseudo label discovery with the aid of fine-grained task tokens $\boldsymbol{\varphi_{i}}$ as mentioned in Sec. 3.2. We directly produce pseudo labels from $\boldsymbol{G_{i}^{\prime}}$: $\boldsymbol{\tilde{Y_{i}}}=\operatorname{Argmax}\left(\operatorname{Softmax}\left(\boldsymbol{G_{i}^{\prime}}\right)\right)$, along with binary masks to select high confidence pixel pseudo labels: $\boldsymbol{M_{i}^{p}}=\operatorname{Max}\left(\operatorname{Softmax}\left(\boldsymbol{G_{i}^{\prime}}\right)\right)>\tau_{i}$, where $\tau_{i}$ is a threshold used to produce binary masks. The loss for pseudo label supervision can be written as:
(4) $\mathcal{L}_{p}=\frac{1}{T}\sum_{i=1}^{T}\left((1-\alpha_{i})\boldsymbol{M_{i}^{p}}\odot\operatorname{L_{i}}\left(\boldsymbol{\hat{Y_{i}}},\boldsymbol{\tilde{Y_{i}}}\right)\right),$
Finally, we sum all of the losses in Eq 2,  3 and  4 to supervise all task features and predictions. The overall losses to optimize the model can be described as: $\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{p}+\mathcal{L}_{f}$, each item is combined with the weight 111 to form the total loss.
4. Experiment
4.1. Experimental Setup
PASCAL-Context. PASCAL-Context (Everingham et al., 2009) contains 4998 and 5105 images for training and testing respectively, which also have pixel-level annotations for semantic segmentation, human-parts segmentation and semantic edge detection. Additionally, we also consider surface normal estimation and saliency detection distilled by (Maninis et al., 2019). We use Adam optimizer with learning rate $2\times 10^{-5}$, and weight decay $1\times 10^{-6}$, and train for 100100100 epochs with batch size 666. We update the learning rate with polynomial strategy and $\gamma=0.9$ for the power factor.
NYUD-v2. NYUD-v2 (Silberman et al., 2012) contains 795 and 654 RGB-D indoor scene images for training and testing respectively. We use the 13-class semantic annotations which is defined in (Couprie et al., 2013), the truth depth annotations recorded by Microsoft Kinect depth camera, and surface normal annotations which are produced in (Eigen and Fergus, 2015). Following the setting in (Liu et al., 2019; Li et al., 2022), we use 288×384288$\times$ 384288 × 384 image resolution to speed up training. We use Adam optimizer with a learning rate of $1\times 10^{-4}$, and train all models for 400 epochs with batch size 888. We update the learning rate every 100100100 epoch with $\gamma=0.5$ as the multiplying factor.
Cityscapes. Cityscapes (Cordts et al., 2016) contains 2975 and 500 street-view images for training and testing respectively. We used the projected 7-class semantic annotations from (Liu et al., 2019), and disparity maps to serve as depth annotations. Following the setting in (Liu et al., 2019; Li et al., 2022), we use 128×256128$\times$ 256128 × 256 image resolution to speed up training. The optimizer and learning rate scheduler are set as the same as NYUD-v2.
Model Setting. Following (Li et al., 2022), we use SegNet (Badrinarayanan et al., 2017) for NYUD-v2 and Cityscapes, ResNet-18 (He et al., 2016) for PASCAL-Context as the backbone of our single task learning (STL) baselines, and the multi-task baseline (MTL) is built from it, which consists of a shared backbone encoder and several task-specific decoding heads. For the learning of HiTTs, we follow (Brüggemann et al., 2021; Li et al., 2018; Zeisl et al., 2014), and perform a discrete quantization of the label space of continuous regression tasks such as Depth. and Normal. This discrete quantization does not contribute to multi-task learning performance as analyzed in (Brüggemann et al., 2021), so we ensure a fair comparison with other works.
Data Preparation. We follow the setting of (Ye and Xu, 2024) to process PASCAL-Context, and the setting of (Li et al., 2022; Liu et al., 2019; Nishi et al., 2024) to process NYUDv2 and Cityscapes, and form two partially annotated settings (Li et al., 2022): (i) one-label: for each input image, it is only associated with one task annotation; (ii) random-labels: each image has at least one and at most $N-1$ tasks with corresponding task annotations, in the set of $N$ tasks. Additionally, we provide two extra settings full-labels and few-shot in the supplementary material to further validate the effectiveness of our method.
Training Pipeline. We first train the multi-task model with HiTTs on all labeled task data. Then the weights of the network and tokens are fixed, and used to produce hierarchical supervision on both feature and prediction spaces. We produce the pseudo label in an offline manner according to (Xie et al., 2020), which is labeling on clean image without data augmentation, and training on augmented images and pseudo labels to enforce consistent predictions. In (Xie et al., 2020), this method is only applied to classification tasks while we extend the utilization to general dense prediction tasks. After the pseudo labels are saved, we use them along with the ground-truth labels to jointly train the multi-task model from scratch. Additionally, we also use the pseudo feature supervision produced by this pretrained multi-task model for feature regularization during the optimization process. Both hierarchies of the discovered supervision signals ensure that all task predictions will obtain pixel-wise supervision for multi-task optimization to gain better generalization ability on unlabeled data.
Evaluation Metrics. We use multiple metrics for each task to evaluate the performance. The metrics include: mIoU (mean intersection over union), AbS / AbR (absolute error / absolute-relative error), maxF (maximal F-measure), mErr (mean of angle error), odsF (optimal dataset scale F-measure). Additionally, to better evaluate the proposed method, we also use $\Delta_{MTL}$ from (Vandenhende et al., 2021) to evaluate the overall improvement of the multi-task performances of all the tasks.
Figure 5. Comparisons with SoTA works on NYUDv2. Ours shows both clear semantic boundaries and accurate geometry estimations, indicating the effectiveness of cross-task feature-token interactions.
Table 1. Quantitative comparison on PASCAL-Context under the one-label and random-labels setting. (P) and (F) represent the Prediction Diffusion and Feature Diffusion modes for DiffusionMTL (Ye and Xu, 2024), and MTDNet is the Multi-Task Denoising Diffusion Networkt (Ye and Xu, 2024). The Mapping Network is the extra encoder used for task mappings in (Li et al., 2022; Nishi et al., 2024). “*” denotes the re-implemented results from (Ye and Xu, 2024). Our method performance outperforms previous methods while using significantly fewer model parameters.
# labels Method MTDNet Mapping #Params FLOPS Semseg Parsing Saliency Normal Boundary MTL Perf
[…]
One-Label
STL 219M 817G 50.34 59.05 77.43 16.59 64.40 -
MTL baseline 157M 608G 49.71 56.00 74.50 16.85 62.80 -2.85
SS (Li et al., 2022) - - 45.00 54.00 61.70 16.90 62.40 -
XTC (Li et al., 2022) - - 49.50 55.80 61.70 17.00 65.10 -
XTC* (Li et al., 2022) 173M 608G 55.08 56.72 77.06 16.93 63.70 0.37
JTR* (Nishi et al., 2024) 173M 608G 50.29 54.78 78.35 17.97 63.66 -3.12
DiffusionMTL (P) (Ye and Xu, 2024) 133M 628G 59.43 56.79 77.57 16.20 64.00 3.23
DiffusionMTL (F) (Ye and Xu, 2024) 133M 676G 57.78 58.98 77.82 16.11 64.50 3.65
Ours 62M 493G 61.24 57.52 78.35 15.75 67.70 6.09
Random-Labels
STL 219M 817G 51.51 57.90 80.30 15.24 67.80 -
MTL baseline 157M 608G 62.23 55.88 78.67 15.47 66.70 2.44
SS (Li et al., 2022) - - 59.00 55.80 64.00 15.90 66.90 -
XTC (Li et al., 2022) - - 59.00 55.60 64.00 15.90 67.80 -
XTC* (Li et al., 2022) 173M 608G 62.44 55.81 78.56 15.45 66.80 2.52
JTR* (Nishi et al., 2024) 173M 608G 57.21 53.18 79.98 16.48 66.20 -1.60
DiffusionMTL (P) (Ye and Xu, 2024) 133M 628G 63.68 55.84 79.87 15.38 66.80 3.44
DiffusionMTL (F) (Ye and Xu, 2024) 133M 676G 62.55 56.84 80.44 14.85 67.10 4.27
Ours 62M 493G 65.19 56.35 81.70 14.80 67.90 5.73
Figure 6. Comparisons of task score maps produced by global task tokens and fine-grained task tokens on Cityscapes.
Table 2. Comparison on NYUD-v2 under one-label and random-labels settings.
Setting Model Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU AbS mErr (%)↑
One-Label
STL 29.28 0.7182 30.1971 -
MTL baseline 30.92 0.5982 31.8509 5.61
MTAN (Liu et al., 2019) 30.92 0.6196 30.0278 6.63
XTC (Li et al., 2022) 33.46 0.5728 31.1492 10.46
JTR (Nishi et al., 2024) 31.96 0.5919 30.8000 8.25
Ours 35.81 0.5540 28.5131 16.91
Random-Labels
STL 34.49 0.6272 27.9681 -
MTL baseline 35.49 0.5503 29.9541 2.69
MTAN (Liu et al., 2019) 35.96 0.6120 28.6933 1.36
XTC (Li et al., 2022) 38.11 0.5387 29.6549 6.19
JTR (Nishi et al., 2024) 37.08 0.5541 29.4400 4.63
Ours 41.78 0.5177 27.3488 13.60
Table 3. Comparison on Cityscapes under one-label setting. “*” denotes the re-implemented results to align the settings.
Setting Model Semseg. Depth. $\Delta_{MTL}$
mIoU AbS (%)↑
One-Label STL 69.69 0.0142 -
MTL baseline 69.94 0.0159 -5.81
MTAN (Liu et al., 2019) 71.12 0.0146 -0.38
XTC (Li et al., 2022) 73.23 0.0159 -3.45
JTR (Nishi et al., 2024) 72.33 0.0163 -5.50
DiffusionMTL (F)* (Ye and Xu, 2024) 73.19 0.0138 3.92
Ours 73.65 0.0135 5.31
4.2. State-of-the-art Comparison
[…]
Comparison on NYUD-v2. We compare our method with (Li et al., 2022; Liu et al., 2019) on NYUD-v2 under both the one-label and random-labels settings, and the quantitative results are shown in Table 2. XTC (Li et al., 2022) is the first work designed for partially annotated multi-task dense prediction. MTAN (Li et al., 2022) is an attention-based MTL network designed for the fully supervised setting, and we train it with our setup. The quantitative results show that our method surpasses them by a large margin on all the metrics of the three tasks. More specifically, ours achieves $+6.45\%~\Delta_{MTL}$ and $+7.41\%~\Delta_{MTL}$ compared with (Li et al., 2022) under the two partial-label settings, respectively. The qualitative comparison with the state-of-the-art method XTC (Li et al., 2022) as shown in Fig. 5 can also confirm the superior performance of our method.
[…]
Table 4. Investigate the effectiveness of different components on NYUD-v2 testing set under one-label setting. $\mathcal{L}_{p}^{*}$ represents the navie pseudo label loss in prediction space without utilizing HiTTs.
Method Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU AbS mErr (%)↑
STL 29.28 0.7182 30.1971 -
MTL baseline 30.92 0.5982 31.8509 5.61
HiTTs w/o. OE 27.38 0.6049 30.5904 2.66
HiTTs w/o. Inter-Task Attention 31.26 0.5966 30.2911 7.79
HiTTs w/o. Intra-Task Attention 31.44 0.5910 30.1432 8.42
HiTTs w/o. $\boldsymbol{\theta_{i}}$ 30.03 0.5823 30.0005 7.38
HiTTs w/o. $\boldsymbol{\varphi_{i}}$ 30.53 0.5842 30.0891 7.76
HiTTs 32.48 0.5844 30.0847 9.98
STL w. $\mathcal{L}_{p}^{*}$ 30.78 0.6693 30.2420 3.93
MTL w. $\mathcal{L}_{p}^{*}$ 33.59 0.5882 29.8174 11.36
HiTTs w. $\mathcal{L}_{f}$ 33.24 0.5708 29.2227 12.88
HiTTs w. $\mathcal{L}_{p}$ 35.22 0.5613 28.8852 15.49
HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ (full method) 35.81 0.5540 28.5131 16.91
Figure 7. Comparison of the training and testing performance on each task with and without hierarchical Pseudo Supervision (P). The model trained with pseudo supervision converges faster on both train and test splits, and gains better performance.
Table 5. Investigate the performance on labeled and unlabeled data on NYUD-v2 training set under the one-label setting. GT and pseudo represents the ground-truth and the pseudo supervision, respectively. Our method clearly shows effective learning on unlabeled training data.
Method Supervision Semseg. Depth. Normal.
GT Pseudo mIoU AbS mErr
MTL baseline 89.00 0.2041 25.9280
34.31 0.5823 31.7697
HiTTs 86.04 0.3016 21.7911
34.69 0.5699 29.8920
HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ 86.63 0.3173 20.9220
37.25 0.5563 28.5169
4.3. Model Analysis
Components of Hierarchical Task Tokens. As shown in Table 4, under the one-label setting on NYUD-v2, we give an ablation study of HiTTs’ key components. Generally, we analyze the role of global task tokens $\boldsymbol{\theta_{i}}$ and fine-grained tokens $\boldsymbol{\varphi_{i}}$, and core designs for token learning process, including the orthogonal embeddings (OE) (3.2), inter- and intra-task attention (3.1). The quantitative results clearly show that both hierarchies of tokens contribute to the multi-task performance and HiTTs boost the model performance by overall $+4.37\%~\Delta_{MTL}$ on all tasks compared with baseline. For the learning process of HiTTs, the inter- and intra-task attention both contribute to the learning process, and the orthogonal embeddings (OE) are essential for generating representative fine-grained task tokens, and without OE, the performance will significantly drop ($-7.32\%~\Delta_{MTL}$).
Effect of Hierarchical Feature Supervision and Label Discovery. To validate that our token-based label discovery is superior to the naive pseudo-labeling process, we compare $\mathcal{L}_{p}$ imposed on different models, including STL, MTL baselines, and ours. The HiTTs perform i) effective cross-task feature-token interactions; ii) consistent label discovery in both feature and prediction space, which yields better pseudo label quality, and significantly surpasses simply applying $\mathcal{L}_{p}$ to STL or MTL baselines without HiTTs. We also analyze the contributions from two types of pseudo supervision losses ($\mathcal{L}_{p}$ and $\mathcal{L}_{f}$) on different hierarchies. As shown in Table 4, both methods boost multi-task performance: $+7.27\%~\Delta_{MTL}$ for $\mathcal{L}_{f}$ and $+9.88\%~\Delta_{MTL}$ for $\mathcal{L}_{p}$ compared with MTL baseline. The combination of both methods achieves better performance ($+11.30\%~\Delta_{MTL}$) than applying them separately, which validates the importance of consistently discovering supervision signals in both hierarchies.
Visualization Results. We visualize: i) The visualization of score maps produced by $\boldsymbol{\theta_{i}}$ and $\boldsymbol{\varphi_{i}}$ in Fig. 6. The visualizations reveal that score maps from global tokens provide a coarse, noisy overview and are biased towards common categories (e.g., focusing only on ”road” in Cityscapes). In contrast, maps from fine-grained tokens are detailed, less noisy, and can identify smaller, less frequent objects (e.g., ”vehicles” and ”pedestrians”). This confirms our hierarchical structure is essential for learning representations at different levels of granularity. ii) The learning curves of metrics on every task in Fig. 7, both training and testing performance are boosted consistently on all tasks with the discovered pseudo supervision (P) on both hierarchies.
Learning Effect on Unlabeled Data. We also study the performance of our method on the labeled and unlabeled data separately on NYUD-v2 training set under one-label setting. As shown in Table 5, for data without labels, the model with HiTTs generalizes better on them, especially on Depth. and Normal, and adding hierarchical supervision will more significantly boost the performance on unlabeled data.
5. Conclusion
In this work, we propose to learn Hierarchical Task Tokens (HiTTs) for both pseudo feature supervision and label discovery under Multi-Task Partially Supervised Learning. The global task tokens are exploited for feature-token cross-task interactions and provide feature-level supervision, while the fine-grained tokens inherit knowledge from global tokens and excavate pixel pseudo labels. Extensive experimental results on partially annotated multi-task dense prediction benchmarks validate the effectiveness of our method.
Supplementary
In this supplementary document, we present: (i) more details about the structure of hierarchical task tokens and training pipeline, (ii) a more comprehensive explanation of experimental implementations, (iii) more quantitative and qualitative experimental results.
Appendix A Model Details
A.1. Training Pipeline
We will discuss utilizing the discovered supervision signals by Hierarchical Task Tokens (HiTTs) for the training process in this section. We first train the multi-task model with HiTTs on all labeled task data. Then the weights of the network are fixed, and used to produce hierarchical supervision on both feature and prediction spaces. We produce the pseudo label in an offline manner according to (Xie et al., 2020), which is labeling on clean image without data augmentation, and training on augmented images and pseudo labels to enforce consistent predictions. In (Xie et al., 2020), this method is only applied to classification tasks while we extend the utilization to general dense prediction tasks. After the pseudo labels are produced, we use them along with the ground-truth labels to jointly train the multi-task model from scratch. Additionally, we also use the pseudo feature supervision produced by this pre-trained multi-task model for feature regularization during the optimization process. Both hierarchies of the discovered supervision signals ensure that all task predictions will obtain pixel-wise supervision for multi-task optimization to gain better generalization ability on unlabeled data.
A.2. Global Task Token Learning
In this section we are going to introduce the inter- and intra-task attention in detail. Given an RGB input $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a multi-task dense prediction framework firstly produces a task-generic representation $\boldsymbol{F_{s}}\in\mathbb{R}^{C\times h\times w}$ through a shared encoder. Considering we have $T$ tasks, and we target decoding task features $\{\boldsymbol{F_{1}},\boldsymbol{F_{2}},\cdots,\boldsymbol{F_{T}}\}$ from $\boldsymbol{F_{s}}$ as well as learning representative global task tokens $\{\boldsymbol{\theta_{1}},\boldsymbol{\theta_{2}},\cdots,\boldsymbol{\theta_{T}}\}$ for each task.
As shown in Fig. 8: i) Inter-Task Learning, which aims to learn explicit global cross-task token affinities $\boldsymbol{A}$, and conduct cross-task interaction accordingly for the global token learning process. ii) Intra-task Learning, which learns task-specific information by globally conducting self-attention between task feature and token pairs.
Firstly, we flatten the shared feature $\boldsymbol{F_{s}}$ into feature tokens with shape $\mathbb{R}^{C\times(hw)}$, and use each global task token to query the shared feature to obtain each task feature $\boldsymbol{F_{i}}$ accordingly. Then, to conduct inter-task attention, all global task tokens are used to produce cross-task affinities that explicitly guide the learning process. The cross-task affinity map $\boldsymbol{A}\in\mathbb{R}^{T\times T}$ is calculated as:
(5) $\boldsymbol{A}=\operatorname{Softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top}),$
where $\boldsymbol{Q}$ and $\boldsymbol{K}$ are individual linear projection of concatenated global tokens $\boldsymbol{\Theta}=[\boldsymbol{\theta_{1}};\boldsymbol{\theta_{2}};\cdots;\boldsymbol{\theta_{T}}]^{\top}$, in which [][$\cdot$][ ⋅ ] indicates the concatenation. After affinity matrix $\boldsymbol{A}$ is calculated, it is used to conduct affine combinations of task features and global task tokens respectively:
[…]
Afterward, the updated tokens and features with cross-task information are involved in the intra-task attention, where we first concatenate every corresponding task token and feature, and perform self-attention on the spatial dimension among each token-feature pair $\left[\boldsymbol{\theta_{i}^{\prime}};\boldsymbol{F_{i}^{\prime}}\right]\in\mathbb{R}^{C\times(hw+1)}$. The global task tokens will further learn more specific and discriminative task representations during this process, and representative task tokens will in turn enhance the feature quality as well. Followingly, we discover pseudo feature supervision with the aid of well-learned global task tokens $\boldsymbol{\theta_{i}^{\prime}}$.
Additionally, for multi-scale backbone features, directly fusing them ignores the various granularity of task representations maintained at different scales. Thus, for multi-scale image backbone, we further propose Multi-scale Global Task Token Learning in order to learn comprehensive multi-scale task relations. The proposed method involves inter-task attention separately at each scale, and then the multi-scale features and tokens are fused before intra-task attention. In this way, the global task tokens gain richer cross-task relations at different scales and are able to maintain stronger representations. We will illustrate this part in detail in Sec. A.4.
Figure 8. Illustration of detail designs of: (i) Inter-Task Attention: predicting cross-task token affinities $\boldsymbol{A}$ from the global task tokens. (ii) Intra-Task Attention: conducting self-attention between task features and tokens.
A.3. Discrete Quantization and Task Losses
In Sec. 3 of the body part, we have discussed how to learn HiTTs. For continuous regression tasks, such as Depth and Normal, we first need to perform a discrete quantization of the label space to provide discriminative supervision for tokens. The goal for quantization is to assign meaningful category bins to each fine-grained token for classification. As analyzed in (Brüggemann et al., 2021), this quantization only changes the way of predicting regression task, but does not contribute to the learning performance. For depth estimation, we follow the setting in (Brüggemann et al., 2021; Li et al., 2018), and divide the range of depth values into several logarithmic bins. Our predicted task logits score $\boldsymbol{G_{i}^{\prime}}$ is used to calculate the soft-weighted sum with each bin and produce final task predictions accordingly. For surface normal estimation, we follow (Brüggemann et al., 2021; Zeisl et al., 2014) and use K-means to learn several unit normal vectors, which serve as clustering centers, and they are also used to generate predictions with $\boldsymbol{G_{i}^{\prime}}$. This process can be expressed as:
(7) $\boldsymbol{\hat{Y_{i}}}=\operatorname{Sum}\left(\boldsymbol{c_{i}}^{\top}\times\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$
where $\boldsymbol{c_{i}}\in\mathbb{R}^{C_{p}}$ represents the center of each bin. In our experiments, the numbers of depth bins and normal cluster centers on NYUD-v2 are 303030 and 202020, respectively. For Cityscapes, we consider 100100100 depth bins as the Cityscapes dataset is captured from outdoor scenarios and have more significant changes in depth. For PASCAL-Context, we select 404040 different unit verctors uniformly from the space to serve as normal cluster centers.
To supervise the task predictions, we can directly impose regression losses on $\boldsymbol{\hat{Y_{i}}}$ for Depth. and Normal:
(8) $\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right),$
where $\operatorname{L_{i}^{reg}}(\cdot)$ can be $\operatorname{L_{1}}$ loss or Angle loss for Depth. or Normal. respectively. Furthermore, in order to gain more discriminative task category information for each token, we also impose classification loss on $\boldsymbol{G_{i}^{\prime}}$. We first extract the corresponding one-hot label from $\boldsymbol{Y_{i}}$ by $\boldsymbol{c_{i}}$, denoting as $\boldsymbol{Y_{i}^{oh}}$, and the loss can be written as:
(9) $\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}},{\boldsymbol{Y_{i}^{oh}}}\right)=-\boldsymbol{Y_{i}^{oh}}\cdot\operatorname{log}\left(\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$
and the overall mixture loss for each task can be written as:
(10) $\operatorname{L_{i}}(\cdot)=\lambda_{i}\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right)+\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}^{\prime}},{\boldsymbol{Y_{i}^{oh}}}\right),$
where $\lambda_{i}=0.1$ if task $i$ is a regression task itself, otherwise $\lambda_{i}=0$.
A.4. Multi-scale Global Task Token Learning
For the implementations on NYUD-v2 and Cityscapes, we use the SegNet (Badrinarayanan et al., 2017) to serve as the shared backbone, which produces single-scale shared features for per-task decoding. However, on PASCAL-Context, we use ResNet-18 (He et al., 2016) as the shared backbone, which can produce multi-scale shared features for decoding. Directly fusing them ignores the various task information maintained in different scales. Thus, we propose to learn the global task tokens on different scales in order to learn more comprehensive task relations.
As shown in Fig. 9, compared with the single-scale global task token learning process, the multi-scale process involves inter-task learning separately on each scale, and then the multi-scale features and tokens are fused before intra-task learning. The multi-scale backbone features $\left\{\boldsymbol{F_{s}}^{(j)}\right\},j=0,1,2,3$ are first flattened on each scale, and each global task token $\boldsymbol{\theta_{i}}$ is projected by a linear layer to produce multi-scale tokens:
(11) $\boldsymbol{\theta_{i}}^{(j)}=\boldsymbol{W^{s\to m}_{i}}^{(j)}\times\boldsymbol{\theta_{i}},\quad j=0,1,2,3,$
where $\boldsymbol{W^{s\to m}_{i}}^{(j)}$ Then, for every feature $\boldsymbol{F_{s}}^{(j)}$ and token $\boldsymbol{\theta_{i}}^{(j)}$ on scale $j$ and task $i$, we query $\boldsymbol{F_{s}}^{(j)}$ to obtain task features $\left\{\boldsymbol{F_{i}}^{(j)}\right\},i=1,2,\cdots,T;j=0,1,2,3$. After that, on each scale, we concatenate features and tokens from every task for Inter-task Learning similar to Sec. 3.1:
[…]
(13) $\boldsymbol{\Theta}^{(j)}$ $=\left[\boldsymbol{\theta_{1}}^{(j)};\boldsymbol{\theta_{2}}^{(j)};\cdots;\boldsymbol{\theta_{T}}^{(j)}\right]^{\top}.$
[…]
(14) $\boldsymbol{F_{i}^{\prime}}$ $=\operatorname{Conv_{i}^{1\times 1}}\left(\left[\boldsymbol{F_{i}^{\prime}}^{(0)};\boldsymbol{F_{i}^{\prime}}^{(1)};\boldsymbol{F_{i}^{\prime}}^{(2)};\boldsymbol{F_{i}^{\prime}}^{(3)}\right]\right),$
(15) $\boldsymbol{\theta_{i}^{\prime}}$ $=\sum_{j=0}^{3}\left(\boldsymbol{W^{m\to s}_{i}}^{(j)}\times\boldsymbol{\theta_{i}^{\prime}}^{(j)}\right).$
Finally, on each task $i$, the updated task features $\boldsymbol{F_{i}^{\prime}}$ and global task tokens $\boldsymbol{\theta_{i}^{\prime}}$ are used for Intra-task Learning. The process is the same as Sec. 3.1.
In this way, we achieve learning global task tokens on multi-scale task features, which gains richer cross-task relations on different scales and maintains stronger representations in the global task tokens.
Figure 9. Illustrations of Single-scale Global Task Token Learning (a) and Multi-scale Global Task Token Learning (b). For the multi-scale features produced by the shared backbone, we use linear layers to produce corresponding task tokens for each scale. Then, in each scale, we query task features and conduct inter-task attention to gain multi-task multi-scale representations. The multi-scale features and tokens are fused before intra-task attention to transfer cross-scale information.
A.5. Broader Applicability and Future Directions
The hierarchical design of our HiTTs is highly generalizable and not limited to the currently studied tasks. Its principles can be readily extended to other fundamental computer vision tasks, such as object detection (Bochkovskiy et al., 2020; Zhu et al., 2020; Cai et al., 2023; Ren et al., 2024; Wang et al., 2022a) and instance segmentation (Chen et al., 2023; Cheng et al., 2022; Schult et al., 2022; Gu et al., 2024), which fundamentally rely on rich, multi-scale pixel representations. The experiment on the PASCAL-Context dataset partially demonstrates this potential, where HiTTs effectively learns from both coarse-grained semantic labels and fine-grained human part annotations simultaneously, showcasing its robust joint-learning capabilities across different granularities.
Furthermore, the core concepts of HiTTs extend beyond the visual domain. Its ability to build cross-task relations upon multi-level representations makes it a promising approach for tasks like hierarchical text classification (Sun and Lim, 2001; Zhou et al., 2020; Stein et al., 2019; Wang et al., 2024, 2022b) or 3D scene-level parsing (Peng et al., 2023; Wang et al., 2025a; Jaritz et al., 2019). Moreover, the unified token-feature interaction mechanism, based on the Transformer architecture, is inherently compatible with diverse data types, paving the way for future exploration in multi-modal learning environments (Lin et al., 2023; Liu et al., 2023; Bai et al., 2023; Wang et al., 2023, 2025b). This adaptability underscores the broad potential of our hierarchical token-based approach for complex multi-task and multi-modal problems.
Appendix B Implementation Details
B.1. Dataset
PASCAL-Context. PASCAL-Context (Everingham et al., 2009) contains 4998 and 5105 images for training and testing respectively, which also have pixel-level annotations for semantic segmentation, human-parts segmentation and semantic edge detection. Additionally, we also consider surface normal estimation and saliency detection distilled by (Maninis et al., 2019). We use Adam optimizer with learning rate $2\times 10^{-5}$, and weight decay $1\times 10^{-6}$, and train for 100100100 epochs with batch size 666. We update the learning rate with polynomial strategy and $\gamma=0.9$ for the power factor.
NYUD-v2. NYUD-v2 (Silberman et al., 2012) contains 795 and 654 RGB-D indoor scene images for training and testing respectively. We use the 13-class semantic annotations which is defined in (Couprie et al., 2013), the truth depth annotations recorded by Microsoft Kinect depth camera, and surface normal annotations which are produced in (Eigen and Fergus, 2015). Following the setting in (Liu et al., 2019; Li et al., 2022), we use 288×384288$\times$ 384288 × 384 image resolution to speed up training. We use Adam optimizer with a learning rate of $1\times 10^{-4}$, and train for 400 epochs with batch size 888. We update the learning rate every 100100100 epoch with $\gamma=0.5$ as the multiplying factor.
Cityscapes. Cityscapes (Cordts et al., 2016) contains 2975 and 500 street-view images for training and testing respectively. We used the projected 7-class semantic annotations from (Liu et al., 2019), and disparity maps to serve as depth annotations. Following the setting in (Liu et al., 2019; Li et al., 2022), we use 128×256128$\times$ 256128 × 256 image resolution to speed up training. The optimizer and learning rate scheduler are set as the same as NYUD-v2.
B.2. Data Preprocessing
We follow the setting of (Li et al., 2022; Liu et al., 2019; Vandenhende et al., 2021; Ye and Xu, 2024) to process training data. For NYUD-v2 and Cityscapes, we use random scaling, cropping and horizon flipping for data augmentation following (Li et al., 2022; Liu et al., 2019). For PASCAL-Context, we follow (Ye and Xu, 2024) and use random scaling, cropping, horizon flipping and photometric distortion for data augmentation. Apart from the two partial settings (one-label and random-labels) we mentioned in Sec.4.1, we also consider two extra settings: (iii) full-labels: each image has labels on every task; (iv) few-shot: one task has only very few labels while other tasks are fully supervised. Both of the extra settings will furthermore show the effectiveness of our method.
B.3. Model Setting
We use SegNet (Badrinarayanan et al., 2017) as the image backbone for our experiments on NYUD-v2 and Cityscapes; and we use ResNet-18 as the image backbone, Atrous Spatial Pyramid Pooling (ASPP) (Chen et al., 2018b) as the task-specific decoding heads. For the threshold $\tau_{i}$ which selects binary mask $\boldsymbol{M_{i}^{p}}$, we use {Semseg: 0.9, Depth: 0.45, Normal: 0.6 } for NYUD-v2 one-label setting, {Semseg: 0.9, Depth: 0.7, Normal: 0.7} for NYUD-v2 random-labels setting. {Semseg: 0.9, Depth: 0.5} for Cityscapes one-label setting, and {Semseg: 0.9, Parsing: 0.85, Normal: 0.7, Sal: 0.7, Edge: 0.9} for PASCAL-Context one-label and random-labels settings.
Table 6. Comparison on NYUD-v2 under one-label and random-labels settings. Our method shows clear performance gain over three tasks, which is consistent with the visualization results.
Setting Model Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU pAcc AbS rmse mErr rmse $\eta_{1}$ $\eta_{2}$ $\eta_{3}$ (%)↑
One-Label
STL 29.28 55.41 0.7182 1.0151 30.1971 37.7115 23.1532 46.4046 58.5216 -
MTL baseline 30.92 58.23 0.5982 0.8544 31.8509 38.6313 19.7083 41.2614 53.6381 0.11
MTAN (Liu et al., 2019) 30.92 57.14 0.6196 0.8477 30.0278 36.7808 21.4199 44.7805 57.5720 3.26
XTC (Li et al., 2022) 33.46 60.95 0.5728 0.8056 31.1492 37.8211 19.8410 42.2268 54.9997 3.60
Ours 35.81 63.22 0.5540 0.7939 28.5131 36.1738 26.4985 50.2357 61.8343 13.23
Random-Labels
STL 34.49 60.52 0.6272 0.8824 27.9681 34.9293 24.6011 49.7888 62.4425 -
MTL baseline 35.49 61.81 0.5503 0.7874 29.9541 36.7726 21.6933 45.0412 57.7516 -1.47
MTAN (Liu et al., 2019) 35.96 61.64 0.6120 0.8272 28.6933 35.3528 23.0253 47.2287 60.1113 -0.48
XTC (Li et al., 2022) 38.11 64.37 0.5387 0.7755 29.6549 36.3992 21.7058 45.4801 58.4236 0.66
Ours 41.78 66.50 0.5177 0.7472 27.3488 34.6820 27.1619 51.8924 63.7670 9.28
Table 7. Comparison on NYUD-v2 under full-labels settings. Our method achieves significantly better performance compared with SoTA multi-task learning works on all of the three tasks.
Model Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU AbS mErr (%)↑
STL 37.45 0.6079 25.94 -
MTL baseline 36.95 0.5510 29.51 -1.91
MTAN (Liu et al., 2019) 39.39 0.5696 28.89 0.03
X-Task (Zamir et al., 2020) 38.91 0.5342 29.94 0.20
Uncertainty (Kendall et al., 2018) 36.46 0.5376 27.58 0.87
GradNorm (Chen et al., 2018a) 37.19 0.5775 28.51 -1.87
MGDA (Désidéri, 2012) 38.65 0.5572 28.89 0.06
DWA (Liu et al., 2019) 36.46 0.5429 29.45 -1.83
XTC (Li et al., 2022) 41.00 0.5148 28.58 4.87
XTC+Uncertainty (Kendall et al., 2018) 41.09 0.5090 26.78 7.58
CCR (Yang et al., 2023b) 43.09 0.4894 27.87 9.04
Ours (HiTTs) 44.32 0.4813 25.76 13.29
[…]
B.4. Evaluation Metrics
We have briefly introduced our evaluation metrics for multiple dense prediction tasks in Sec.4.1. We provide a more detailed description as follows: (i) mIoU: mean intersection over union; (ii) pAcc: per-pixel accuracy; (iii) AbS or AbR: absolute error or absolute-relative error; (iv) rmse: root mean square error (for Normal. we calculate the mean square error of the predicted angles with the ground-truths); (v) mErr: mean of angle error; (vi) odsF: optimal dataset F-measure (Martin et al., 2004); (vii) threshold: for surface normal estimation, we calculate the proportion of pixels with angle error smaller than three thresholds $\eta\in\left\{11.25^{\circ},22.50^{\circ},30^{\circ}\right\}$.
Additionally, to better evaluate the proposed method, we also consider using $\Delta_{MTL}$ proposed by (Vandenhende et al., 2021) to evaluate the overall improvement of the multi-task performances of all the tasks, which is defined as:
(16) $\Delta_{MTL}=\sum_{N}(-1)^{l_{t}}\left(M_{m,t}-M_{s,t}\right)/M_{s,t}$
where $l_{t}=1$ if a lower evaluation value indicates a better performance measurement of $M_{t}$ for task $t$, and $l_{t}=0$ if a higher value is better. Footnote $s$ and $m$ represent the performance of the single-task learning and the multi-task learning respectively. We will show experimental results with all of these metrics to further show the effectiveness of our method.
Table 8. Comparison on Cityscapes under one-label setting.
Setting Model Semseg. Depth. $\Delta_{MTL}$
mIoU pAcc AbS rmse (%)↑
One-Label
STL 69.69 91.91 0.0142 0.0271 -
MTL baseline 69.94 91.62 0.0159 0.0292 -4.92
MTAN (Liu et al., 2019) 71.12 92.35 0.0146 0.0278 -0.72
XTC (Li et al., 2022) 73.23 92.73 0.0159 0.0293 -3.53
Ours 73.65 92.81 0.0135 0.0265 3.45
Table 9. Investigate the effectiveness of different components on NYUD-v2 testing set under one-label setting.
Method Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU pAcc AbS rmse mErr rmse $\eta_{1}$ $\eta_{2}$ $\eta_{3}$ (%)↑
STL 29.28 55.41 0.7182 1.0151 30.1971 37.7115 23.1532 46.4046 58.5216 -
MTL baseline 30.92 58.23 0.5982 0.8544 31.8509 38.6313 19.7083 41.2614 53.6381 0.11
+HiTTs w/o. OE 27.38 55.08 0.6049 0.8626 30.5904 38.1046 23.5233 45.7012 57.1902 2.13
+HiTTs w/o. Inter 31.26 57.99 0.5966 0.8592 30.2911 37.7714 22.7802 46.2590 58.0880 4.51
+HiTTs w/o. Intra 31.44 58.50 0.5910 0.8533 30.1432 37.6719 23.3251 46.4354 58.1508 5.23
+HiTTs w/o. $\boldsymbol{\theta_{i}}$ 30.03 58.21 0.5823 0.8389 30.0005 37.3750 23.4160 46.1824 58.2909 5.08
+HiTTs w/o. $\boldsymbol{\varphi_{i}}$ 30.53 57.18 0.5842 0.8565 30.0891 37.4465 23.2297 45.9603 58.0509 4.60
+HiTTs 32.48 59.61 0.5844 0.8382 30.0847 37.5827 23.9975 46.4790 58.2146 6.51
STL w. $\mathcal{L}_{p}$ 30.78 58.94 0.6693 0.9362 30.2420 37.8601 23.5830 46.4743 58.3739 3.03
MTL w. $\mathcal{L}_{p}$ 33.59 61.79 0.5882 0.8554 29.8174 36.9781 23.2875 45.8803 58.1061 6.89
+HiTTs w. $\mathcal{L}_{f}$ 33.24 60.74 0.5708 0.8200 29.2227 36.9305 25.7968 48.8173 60.2608 9.75
+HiTTs w. $\mathcal{L}_{p}$ 35.22 62.93 0.5613 0.8014 28.8852 36.4316 25.3873 49.1251 60.9806 11.57
+HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ 35.81 63.22 0.5540 0.7939 28.5131 36.1738 26.4985 50.2357 61.8343 13.23
Table 10. Investigate the performance on labeled and unlabeled data of NYUD-v2 training set under one-label setting.
Method Supervision Semseg. Depth. Normal.
GT Pseudo mIoU pAcc AbS rmse mErr rmse $\eta_{1}$ $\eta_{2}$ $\eta_{3}$
MTL baseline 89.00 96.04 0.2041 0.3434 25.9280 32.0824 25.8276 52.0758 65.4135
34.31 61.56 0.5823 0.8375 31.7697 38.5071 19.5042 41.2401 53.8490
HiTTs 86.04 95.00 0.3016 0.4625 21.7911 28.9368 38.8125 63.6280 73.8326
34.69 61.48 0.5699 0.8319 29.8920 37.3331 23.9788 46.4809 58.4659
HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ 86.63 94.89 0.3173 0.4770 20.9220 28.1260 41.2846 66.0712 75.6904
37.25 63.72 0.5563 0.8074 28.5169 36.1536 26.4528 50.0344 61.5778
Table 11. Investigate the cross-task learning effect on NYUD-v2 under the few-shot setting.
  Method   Few-Shot-Semseg   Few-Shot-Depth   Few-Shot-Normal
  mIoU   pAcc   AbS   rmse   mErr   rmse   $\eta_{1}$   $\eta_{2}$   $\eta_{3}$
  STL   5.80   26.06   0.9533   1.2907   47.5281   53.6422   5.4343   17.8888   27.9915
  MTL baseline   16.75   41.01   0.9165   1.2968   40.0456   46.3370   12.0348   26.8520   37.2863
  HiTTs   18.05   44.69   0.7803   1.1272   39.8508   47.1113   14.1108   29.4719   39.5924
  HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$   20.07   45.62   0.7366   1.0206   38.5029   46.9907   17.1082   34.7171   45.1765
Table 12. Comparison of HiTTs with Single-scale (SS) and Multi-scale (MS) Global Task Token Learning on PASCAL-Context under the one-label and random-labels setting.
Setting Model Semseg. Parsing. Norm. Sal. Edge. $\Delta_{MTL}$
mIoU mIoU mErr mIoU odsF (%)↑
One-
Label
STL 47.7 56.2 16.0 61.9 64.0 -
MTL baseline 48.4 55.1 16.0 61.6 66.5 0.59
HiTTs (SS) 51.0 54.7 16.2 61.7 66.1 1.19
HiTTs (MS) 52.3 56.2 15.8 62.0 67.9 3.43
Random-
Labels
STL 60.9 55.3 14.7 64.8 66.8 -
MTL baseline 58.4 55.3 16.0 63.9 67.8 -2.57
HiTTs (SS) 59.1 53.4 15.0 64.1 67.8 -1.60
HiTTs (MS) 60.3 55.3 14.7 64.6 70.2 0.76
B.5. More Quantitative Results
B.5.1. State-of-the-art Comparison
Comparison on NYUD-v2. We compare our method with (Li et al., 2022; Liu et al., 2019) on NYUD-v2 under both the one-label and random-labels settings, and the quantitative results are shown in Table 6. XTC (Li et al., 2022) is the first work designed for partially annotated multi-task dense prediction. MTAN (Li et al., 2022) is an attention-based MTL network designed for the fully supervised setting, and we train it with our setup. The quantitative results show that our method surpasses them by a large margin on all the metrics of the three tasks. More specifically, ours achieves $+9.63\%~\Delta_{MTL}$ and $+8.62\%~\Delta_{MTL}$ compared with (Li et al., 2022) under the two partial-label settings, respectively.
Our HiTTs can also be applied on full-labels setting, which utilizes cross-task relations and task-token interactions to fertilize the multi-task learning process. We compare with some of the recent works, including multi-task interaction works like MTAN (Liu et al., 2019), X-Task (Zamir et al., 2020) and CCR (Yang et al., 2023b); and multi-task loss weighting strategies like Uncertainty (Kendall et al., 2018), GradNorm (Chen et al., 2018a), MGDA (Désidéri, 2012) and DWA (Liu et al., 2019). As shown in Table 7, our method still clearly surpasses all of the SOTA works ($13.29\%~\Delta_{MTL}$ overall), indicating the effective cross-task interaction brought by HiTTs. Additionally, with the aid of feature-level supervision loss $\mathcal{L}_{f}$, which is supported by global task tokens, our method can achieve $14.64\%~\Delta_{MTL}$ overall on the three tasks.
[…]
B.5.2. Model Analysis
Effect of Hierarchical Task Tokens. As shown in Table 9, under the one-label setting on NYUD-v2, we give a more-detailed ablation study on the key components of hierarchical task tokens (HiTTs), including both hierarchies of the task tokens, and the inter-task (Inter) and intra-task (Intra) learning of global task tokens, and orthogonal embeddings (OE) of fine-grained task tokens, and two types of pseudo supervision losses ($\mathcal{L}_{p}$ for pseudo label loss and $\mathcal{L}_{f}$ for feature supervision loss) on different hierarchy. The quantitative results clearly show that every component contributes to the multi-task performance on all the metrics and on all the tasks over the MTL baseline. HiTTs boost the model performance by conducting cross-task interaction and encouraging high-confidence predictions, and leads to an overall $+9.64\%~\Delta_{MTL}$ on all tasks compared with baseline. However, for the learning process of HiTTs, the orthogonal embeddings (OE) are essential for generating representative fine-grained task tokens, and without OE, the performance will significantly drop ($-4.38\%~\Delta_{MTL}$), especially on Semseg. which requires more discriminative category information. The inter-task and intra-task learning processes are also important since without either of them, the learning of cross-task relations and task representations will be affected, resulting in $-2.00\%~\Delta_{MTL}$ and $-1.28\%~\Delta_{MTL}$ performance after removing them respectively from the learning process of HiTTs.
Effect of Hierarchical Feature Supervision and Label Discovery. We analyze the contributions from both feature-level and prediction-level supervision with all of the metrics, as shown in Table 9, both methods boost multi-task performance, and $\mathcal{L}_{p}$ contributes more since fine-grained task tokens contain more specific and discriminative task information and are directly involved in the formation process of task predictions. The combination of both methods achieves better performance than applying them separately, which validates the importance of consistently discovering supervision signals in both hierarchies. Comparing $\mathcal{L}_{p}$ imposed on different models, including STL, MTL baselines, and our HiTTs, our token-based pseudo-label discovery is much better. Since MTL shares a backbone that learns stronger representations on all tasks, MTL produces pseudo labels with better quality and surpasses STL a lot in performance ($+3.86\%~\Delta_{MTL}$). Our HiTTs perform i) consistent label discovery in both feature and prediction space; ii) effective cross-task feature-token interactions, which furthermore enhance the quality of pseudo labels, and bring extra $+4.68\%~\Delta_{MTL}$ overall.
We also study the performance of our method on the labeled and unlabeled data separately on NYUD-v2 training set under the one-label setting. As shown in Table 10, for data without labels, the model with HiTTs generalizes better on them, especially on Depth. and Normal, and adding hierarchical supervision will more significantly boost the performance of unlabeled data.
[…]
Analysis of Multi-Scale Global Task Token Learning. As we illustrated in Sec. A.4, we adopt multi-scale global task learning with ResNet-18 backbone on PASCAL-Context. To validate the effectiveness of learning global task tokens on multi-scale features, we compare the performance of Global Task Token Learning with single-scale and multi-scale backbone features respectively in Table 12. As shown in the table, HiTTs with single-scale (SS) Global Token Learning surpass the MTL baseline on both one-label and random-labels settings, with overall $+0.60\%~\Delta_{MTL}$ and $+0.97\%~\Delta_{MTL}$ on all tasks respectively, and the multi-scale (MS) Global Token Learning further enhances the performance to $+2.84\%~\Delta_{MTL}$ and $+3.33\%~\Delta_{MTL}$ on all tasks, which indicates the effectiveness of applying global token learning on multi-scale features.
Analysis of Threshold Hyperparameter $\tau_{i}$. We discuss the effect of the confidence threshold hyperparameter, $\tau_{i}$. This threshold is chosen to ensure that high-confidence pixel predictions are masked out to serve as pseudo-labels. To analyze the sensitivity of our method to this hyperparameter, we conduct an ablation study on the NYUD-v2 dataset under the one-label setting. As shown in Table 13, we vary $\tau_{i}$ across a wide range from 0.30.30.3 to 0.950.950.95. The results demonstrate that the performance on all tasks remains remarkably stable with only minor fluctuations. Furthermore, the overall multi-task learning gain ($\Delta_{MTL}$) is consistently and significantly positive across all tested values. This demonstrates that our method is not sensitive to the choice of $\tau_{i}$, showcasing its robustness.
[…]
Implementations with Different Backbones. To demonstrate that our method can be flexibly implemented on different image backbones, we perform additional experiments implementing our HiTTs with a Vision Transformer (ViT (Dosovitskiy et al., 2020)) backbone. The results on NYUD-v2 are presented in Table 16. The quantitative results clearly show that when equipped with the more powerful ViT-base backbone, our method achieves a substantial performance improvement across all tasks under both the one-label and random-labels settings. For instance, in the random-labels setting, using ViT-base boosts the Semantic Segmentation mIoU from 47.73 to 61.88. This not only confirms the flexibility of our approach but also highlights its potential to achieve even greater performance when paired with more advanced backbone architectures.
Analysis of Incorporating Dense FixMatch (Pieropan et al., 2022) on NYUD-v2. Our method can also incorporate general semi-supervised dense prediction strategies, e.g. Dense FixMatch (Pieropan et al., 2022). The quantitative results on the NYUD-v2 one-label setting are shown in Table 14. The results indicate that Dense FixMatch provides consistent performance improvements. This demonstrates the efficacy of Dense FixMatch as a versatile component for enhancing various dense prediction tasks in a semi-supervised context.
Table 13. Ablation study on the hyperparameter $\tau_{i}$.
$\tau_{i}$ Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU mErr mErr (%)↑
0.95 46.67 0.4585 25.48 3.11
0.90 47.10 0.4573 25.43 3.57
0.70 46.89 0.4621 25.36 3.17
0.50 47.00 0.4690 25.35 2.79
0.30 47.11 0.4706 25.33 2.78
Table 14. Comparisons with incorporating general semi-supervised dense prediction method (Pieropan et al., 2022) on NYUD-v2 one-label setting.
   Method    Semseg.    Depth.    Normal.
   mIoU    mErr    mErr
   Ours    47.30    0.4539    25.40
   Ours+(Pieropan et al., 2022)    47.72    0.4502    25.33
Table 15. Comparison on NYUD-v2 under the settings of DiffusionMTL (Ye and Xu, 2024). (P) and (F) represent the Prediction Diffusion and Feature Diffusion modes for DiffusionMTL (Ye and Xu, 2024). “*” denotes the re-implemented results to align the settings. Our method performance outperforms previous methods.
Setting Model Semseg. Depth. Normal. $\Delta_{m}$
mIoU $\uparrow$ absErr $\downarrow$ mErr $\downarrow$ (%) $\uparrow$
One-Label
STL 45.28 0.4802 25.93 -
MTL baseline 43.92 0.5138 26.44 -3.99
SS (Li et al., 2022) 27.52 0.6499 33.58 -
XTC (Li et al., 2022) 30.36 0.6088 32.08 -
XTC* (Li et al., 2022) 43.97 0.5140 26.30 -3.79
DiffusionMTL (P) (Ye and Xu, 2024) 44.97 0.5137 26.17 -2.86
DiffusionMTL (F) (Ye and Xu, 2024) 44.47 0.5059 25.84 -2.27
Ours* 47.30 0.4539 25.40 3.99
Random-Labels
STL 48.25 0.4792 24.65 -
MTL baseline 45.93 0.4839 25.53 -3.12
SS (Li et al., 2022) 29.50 0.6224 33.31 -
XTC (Li et al., 2022) 34.26 0.5787 31.06 -
XTC* (Li et al., 2022) 46.03 0.4811 25.97 -3.44
DiffusionMTL (P) (Ye and Xu, 2024) 47.44 0.4803 25.26 -1.45
DiffusionMTL (F) (Ye and Xu, 2024) 46.82 0.4743 24.75 -0.77
Ours* 47.73 0.4510 24.86 1.32
Table 16. Ablation study of our method with different backbones on NYUD-v2. The results show that our method is flexible and achieves significant gains when paired with a more powerful ViT backbone.
Setting Backbone Semseg. Depth. Normal.
mIoU $\uparrow$ absErr $\downarrow$ mErr $\downarrow$
One-Label ResNet-18 47.30 0.4539 25.40
ViT-base 58.38 0.3740 23.65
Random-Labels ResNet-18 47.73 0.4510 24.86
ViT-base 61.88 0.3979 23.03
Figure 10. Comparison of the task confidence map before and after refined by the fine-grained task tokens, which greatly encourage high-confidence predictions on all tasks (red color represents high-confidence areas). For noisy data like the second photo taken in a dark environment, this enhancement are more significant.
B.6. More Qualitative Results
We provide more qualitative results mainly from four parts: more visualization of the fine-grained token distributions, more comparisons of task score maps produced by HiTTs, more qualitative prediction comparisons, and the quality of generated pseudo labels.
Role of Fine-grained Task Tokens. To illustrate the role of our fine-grained task tokens, we visualize task confidence maps before and after the refinement process in Fig. 10. The fine-grained tokens significantly enhance prediction confidence across all tasks, as indicated by the expansion of high-confidence areas (represented in red). This enhancement is particularly evident in challenging scenarios, such as the noisy image captured in a dark environment (second row), where the refined map shows a marked improvement in clarity and confidence.
Visualization of token distributions. We also provide visualization analysis to show the distributions of fine-grained task tokens on Cityscapes. In Fig. 11, with the aid of OE, the self-correlation map of tokens will be more diagonal, and the distributions after PCA have better clusters in 3-dimensional feature space.
Comparisons of Task Score Maps. We visualize score maps produced by global task tokens and fine-grained task tokens respectively on each task. As shown in Fig. 12, we conduct visualization on both NYUD-v2 and Cityscapes datasets. The score maps indicate the response of task features to task tokens, and the response patterns of feature maps on different tasks are very different, e.g. Semseg. features highlight areas with distinguish semantics, Depth. features focus on areas with a certain depth range and Normal. features focus on surfaces with the same orientation.
Comparing the score maps produced by tokens from different hierarchies, we find that score maps produced by global task tokens are relatively rough and noisy, while those generated by fine-grained task tokens have finer granularity and less noise, which shows the hierarchy of the HiTTs learning process. Also, we observed that the high-light areas of global task tokens are monotonous, while fine-grained task tokens can highlight more details. This phenomenon is clearly observed in Cityscapes, since the ground truths of this dataset follow the long-tail distribution, thus the global task-tokens tend to learn the category with more pixel samples, and consequently always highlight the road area as shown in Fig. 12. However, the fine-grained tokens can give attention to more details, including the vehicles and pedestrians with fewer pixel samples. Thus, it is necessary to design a hierarchical structure for tokens to learn representations with different granularity.
Qualitative prediction comparisons with SOTA works. We additionally provide comparisons with SOTA works on NYUD-v2 and Cityscapes. As shown in Fig. 13, we compare with MTL baseline, XTC (Li et al., 2022) on NYUD-v2 three tasks, and with MTL baseline, MTAN (Liu et al., 2019), XTC (Li et al., 2022) on Cityscapes two tasks. Our method shows clearly better performance in semantic understanding and accurate geometry estimations (including depth and normal estimation), indicating the effectiveness of our method.
Visualization of Pseudo Labels. In Fig. 14, we show pseudo task label maps generated by fine-grained task tokens. The pseudo label on different tasks has good quality without ground-truth supervision, which proves the effective cross-task learning and strong generalization ability brought by HiTTs.
Figure 11. Visualization of self-affinities heatmap (left) and PCA for distributions (right) of fine-grained task tokens of the two tasks on NYUDv2 and Cityscapes validation sets. With orthogonal embeddings, the affinities between different tokens are low and the clustering of token distributions on each category is better.
Figure 12. Comparisons of task score maps produced by global task tokens and fine-grained task tokens. The upper part is the visualization of samples on NYUD-v2 while the lower part is on Cityscapes.
Figure 13. Comparisons with SOTA works on NYUD-v2 (upper part) and Cityscapes (lower part).
Figure 14. Quantitative analysis of the quality of pseudo labels generated by gloabl task tokens.